# Natural, Artificial, and Human Intelligences

### v1

@@ Abstract
[…]
@@ 1 Statement of the Problem: are Chatbots Intelligent?
We are witnessing one of the most important and transformative technological revolutions in our history, the emergence of human-like intelligence. A discussion with a modern chatbot, such as chat GPT (we refer to large language models as chatbots in a generic way), is undeniably human-like and several studies have been establishing that chatbots display human cognitive traits, including similar biases (Binz and Schulz,, 2023; Suri et al.,, 2024), the recognition of emotional states (Schaaff et al.,, 2023) and empathy (Roshanaei et al.,, 2024), and other mental processing similarities (Mei et al., (2024); Strachan et al., (2024)). One effect of this revolution is existential: we, humans, often assume our place in the world is unique, by virtue of the uniqueness of our intelligence. Is human intelligence unique?
@@
A more practical question is whether chatbots will take our jobs. Together with being human-like, chatbots also encapsulate, in some sense, all human knowledge. Asking GPT-4 whether it will take my lecturer job, it is humble and reassuring: “As a university lecturer, you’re definitely in a role where expertise, creativity, and human connection are key. While AI, like ChatGPT, can help with things like answering questions, explaining concepts, or generating teaching materials, it’s not likely to replace the heart of your job.” However, increasingly in higher education, new policies are introduced to regulate the use of chatbots, given their dual ability to generate plausible coursework and to mark it with some competence (e.g., Gultom et al., (2024)), questioning the role of humans in the process. Moreover, there are already indications that chatbots can carry out research roles (Lehr et al.,, 2024).
@@
The observation that technological advancements can lead to job losses is hardly new — there is a name for this, “technological unemployment” (Peters,, 2020) — and there are some compelling examples, e.g., the reduction of the U.S. workforce employed in agriculture between 1900 (41%) and 2000 (2%, Autor, (2014)). Unsurprisingly, such concerns have been raised for chatbots too (Callum, 2016). The balancing perspective is that technological innovation also leads to new jobs (Autor, (2015); Mincer and Danninger, (2000)) and over-productivity could be compensated in other ways (such as universal income). Additionally, some jobs require a human touch, do not depend on high levels of intelligence, or are even unnecessary (perhaps the result of vested interests). However, there is a unique role of chatbots in this debate in that they are not ‘just’ another technological innovation with potential to eliminate some jobs: in all cases of technological automation, the replacement concerned tasks that were either too laborious (e.g., agricultural tasks) or seen as beyond the competencies which we have considered central to our notion of human intelligence (e.g., multitudes of rote arithmetical operations in accounting). Instead, with chatbots, we are faced with an artificial agent which seems just like us, concerning some competencies which we have considered distinctly human.
@@
Overall, the consideration of whether chatbots can replace humans in various jobs is important and complements the existential question of whether chatbots are human-like. The present focus is intelligence. Can chatbots be considered intelligent? A difficulty in addressing this question is that our understanding of intelligence is itself incomplete. We consider some relevant ideas below. Briefly, on the one hand, it appears that traditional notions of intelligence can be covered by chatbots – and, moreover, a careful discussion reveals challenges to a uniqueness premise for human intelligence. On the other hand, the fundamental question is whether there is a special ingredient in human intelligence, which allows us to rise to apparent amazing intellectual accomplishments. What is it about our minds which, for example, has enabled us to invent quantum mechanics?
@@ 2 What is Intelligence: Psychological Approaches
Much of the psychological tradition regarding intelligence concerns psychometric theories, which aim is to explain individual differences in test scores, on the basis of factors (e.g., see (Sternberg,, 2012)). The earliest proposal of this kind is due to Spearman, (1904), who argued for two factors, a general one (later associated with ‘mental energy’) and a specific one, comprised of competencies for particular tasks. Thurstone, (1938), instead, identified seven ”primary mental abilities” as the determinants of intelligence. These mental abilities include verbal fluency, spatial visualization, and perceptual speed. While originally Spearman, (1904) and Thurstone, (1938) were interpreted as antithetical, this is not necessarily the case. One way to reconcile the two views is if the general vs. more specific intelligence factors are hierarchical, so that the latter informs specific aspects of the former. This hierarchical approach has been taken up by other intelligence psychologists, such as Carroll, (1993), who expanded both the specific and general abilities.
@@
This literature has recognised that some abilities reflect knowledge, while others processing power (in some general sense). Cattell, (1943) introduced the distinction between fluid and crystallized abilities, with the former concerning performance that depends on prior knowledge minimally (e.g., problem-solving on a novel task) and conversely for the latter (e.g., vocabulary tasks). There is a converse relation between these two kinds of abilities and age, with fluid decreasing, while crystallized increasing, with age (e.g., Ghisletta et al., (2012)).
@@
A more processing account of intelligence has been that it is characterized by the speed of information processing (Deary et al.,, 2001), sometimes in simple perceptual tasks, such as whether two lines are the same length (MacKintosh,, 1981) (but see (Kiesling and Hentschel,, 1988)). There have been related proposals; for example, Hunt et al., (1973) considered the speed with which information is ’absorbed’ into memory and subsequently retrieved. Such approaches do not depend on extent of knowledge or other culture-specific skills, such as social skills.
@@
There have been other attempts to understand intelligence in more cognitive terms, e.g. in terms of working memory capacity (Conway and Engle,, 1996), underpinned by high correlations between measures of working memory and performance with reasoning ability (e.g., (Kyllonen and Christal,, 1990; Ackerman et al.,, 2005)). Interestingly, in cognitive science, working memory capacity limitations have been argued to enable learning in complex domains (Newport,, 1988).
@@
Another processing idea is that intelligence relates to the ability to solve problems and achieve goals. Sternberg, (1985) proposed that intelligence is characterized by three abilities: practical intelligence (environment interaction), creative intelligence (new problems and situations), and analytical intelligence (primarily problem-solving). Problem-solving ability has had a prominent role in other proposals (e.g., (Danner et al.,, 2011; Hambrick et al.,, 2019)), as well as often being a component in most others (starting from (Spearman,, 1904)). Sternberg, (1985)’s theory is an example of approaches positing multiple types of intelligence. Gardner, (2006) proposed a longer list, including musical, interpersonal, spatial, and linguistic intelligence, and Salovey and Mayer, (1990) introduced the idea of emotional intelligence as the ability to perceive and utilize emotions in cognition.
@@
Psychometric approaches to intelligence have demonstrated practical value. For example, they predict longevity (Deary,, 2001; Deary et al.,, 2007), health outcomes (Hart et al.,, 2003), job performance (Schmidt and Hunter,, 2004; Ericsson and Pool,, 2016), and academic success (Deary et al.,, 2007).
@@
Particular aspects of intelligence have been prioritised in intelligence testing: the approaches of (Spearman,, 1904), Binet and Simon, (1905) were translated into educational and cultural settings that emphasized the importance of written answers to prearranged test questions (Rogele,, 2020). A preference for written demonstrations of intelligence, particularly in English-speaking educational systems, has profoundly affected the way artificial intelligence has been viewed and pursued, a relationship that we explore throughout the rest of this paper.
@@
Summing up some relevant points, first, regardless of whether a single general intelligence is postulated or not, it is assumed that there are cognitive abilities that correlate with intelligence/the intelligences (Corr and Krupic,, 2024). Second, some of these abilities concern processing characteristics, for example, processing speed or memory efficiency. Third, intelligence involves several aspects, e.g., problem-solving, linguistic/verbal, emotional, social, etc. Regarding the present focus, can the psychometric literature help us understand what, if anything, sets apart human intelligence compared to biological species and artificial agents? We can eliminate some possibilities. For example, regarding non-human animals, they often demonstrate the competencies (at least to a degree) cited for humans too, such as spatial, emotional, musical abilities etc. (cf. Gardner, (2006)). Moreover, artificial agents, including chatbots, exceed human capacity in essentially all processing characteristics (speed of processing, memory, information absorption and retrieval, problem solving ability, at least in specific categories). The linguistic competence of most chatbots is mostly equivalent to that of humans, even if the extent of understanding is unclear (Mahowald et al.,, 2024). Despite the many profound insights in the psychometric literature on intelligence, there is not very much concerning the characteristics (if any) that make it special.
@@ 3 Animal Intelligence
Humans are not the fastest, toughest, most enduring etc. creatures, but we are distinguished for our cultural, technological, and scientific accomplishments – our achievement in these respects surpasses that of all other creatures. This has allowed us to consider ourselves the most intelligent species. This is probably true, but there is a question of whether other species are intelligent like us, but to a lesser extent, or less intelligent in some fundamental way.
@@
There are signs of intelligence throughout the animal kingdom (Bukart et al., 2017). For example, dogs have highly developed social awareness, including sensitivity to human body language, human gestures, and parts of language (Hare et al.,, 2002; Kaminski et al.,, 2004). Dogs are able to solve a variety of logical/ reasoning problems, such as a means end task (Range et al.,, 2011) or spatial puzzles (Barrera et al.,, 2015). Theory of mind is the ability to attribute mental states (such as beliefs and desires) to others. It is typically considered an integral part of intelligence, especially social intelligence (Navarro,, 2022). There have been arguments that dogs and other primates have a theory of mind (Catala et al.,, 2017; Premack and Woodruff,, 1978), though there is controversy (Heyes,, 1998). Chimpanzees have extensive tool-using abilities, including for food gathering/ hunting, grooming, and communication (Koops et al.,, 2015). Chimpanzees can recognize abstract symbols (Rumbaugh and Washburn,, 2003) and can innovate solutions to unfamiliar problems (Bandini and Harrison,, 2020). Similar evidence has accumulated for other species, e.g., corvids, who appear to have problem-solving abilities comparable to those of chimpanzees (Güntürkün et al.,, 2024). Perhaps unsurprisingly, there is much evidence regarding musical intelligence in birds, e.g., relating to rhythm perception and processing (Fitch,, 2015; Patel,, 2006) or the structural properties of birdsong (Kawaji et al.,, 2024; Zhang et al.,, 2023). Further, there is some evidence of musical synchronisation amongst birds, though for duetting birds at most (Hall,, 2009; Mann et al.,, 2005). Overall, bird musical ability does not appear to match the complexity of human music (Lerdahl and Jackendoff,, 1983; Levitin,, 2006).
@@
What about, for example, the homing instinct of a pigeon or the way a leopard calculates the exact time and distance before it pounces onto prey? For example, pigeons have been known to successfully return home across distances exceeding 100km (Schmidt-Koenig and Schlichte,, 1972). We would not call an individual pigeon intelligent for demonstrating its homing instinct, in the same way we would not call an individual human intelligent for learning to walk or ride a bicycle — such skills are considered default competencies that most members of a species are endowed with, as a result of evolutionary adaptation. There is intelligence at the species level, allowing for different competencies, developed through evolution (e.g., (Jerison and Jerison,, 1988)). A similar case can be made regarding the remarkable ingenuity in, for example, the way ants or bees build their nests (Winston,, 1987).
@@
Beyond evolutionary intelligence, for some species, there is collective intelligence, the ability to achieve more than what an individual could accomplish, through cooperation. In the animal kingdom, for example, ants use pheromones to communicate about food sources or dangers and often work together in foraging tasks (Hölldobler and Wilson,, 1990); the way African wild dogs coordinate in hunting has sophisticated characteristics, including role differentiation and vocal communication (Creel and Creel, (2002); see for an example regarding rats Schuster, (2002)). However, no other species has been more successful in transcending what is possible for an individual than humans. Fernbach and Sloman, (2018) coined the term ‘community of knowledge’, for the idea that even seemingly simple aspects of our lives are underwritten by massive knowledge contributions of which we are mostly ignorant. Our ability to form sophisticated communities of knowledge is fundamentally linked to the unique expressivity of human language (e.g., Lakoff and Johnson, (2003); Pinker, (1995)) and the fact that our language has a written form (Latour and Woolgar,, 1987). Written language has enabled global communities of knowledge; arguably, it is the cornerstone of our civilization (Goody, (1977); concerning the interaction between climate, the durability of early writing mediums, such as papyrus, and the emergence of early civilizations see Gaudet, (2014)).
@@
Is there any evidence for art, science, technology amongst non-humans? Regarding art, some animals can learn to discriminate based on the aesthetics of pictures and show preference towards higher aesthetics (e.g., songbirds, for musical stimuli; Watanabe, (2012)). Some birds create and decorate elaborate structures to attract mates (Borgia,, 1985) and some animals can create pictorial patterns in the sand or using rudimentary tools, which resembles artistic expression (Pelé et al.,, 2021). However, such evidence is a far cry from demonstrations that non-human animals show human-like appreciation of these behaviours and their outputs, in the way we do for art (English et al.,, 2014). Regarding science, the scientific endeavour implies experimentation and exploration of testable hypotheses, to identify general principles and so generalize from known situations to novel ones. There is no evidence that animals can do science. However, there is some evidence that animals can experiment and adjust strategies, create artifacts, and develop rudimentary tools. For example, chimpanzees make tools to help them forage for ants or termites, through a process of successive improvement before the desired form is achieved, sometimes before actual tool use (indicating that they have an idea of the required form for a tool prior to use; Boesch and Boesch, (1990)); and New Caledonian crows have been reported to construct hooked tools, to aid them in insect capture (Hunt et al.,, 1973). Moreover, there is evidence of animals adapting behaviour from trial and error and changing circumstances (e.g., Fiorito and Scotto, (1992)). Such evidence indicates that animals are capable of technology and innovation, though it appears that their only path is through gradual trial and error, as opposed to the sudden realization which often accompanies human technological innovation (famous examples are penicillin and the discovery of the Benzene chemical structure, Fleming, (1929); Kekulé, (1865), as cited in (Rocke,, 2015)).

### v2

@@ Abstract
Human achievement, whether in culture, science, or technology, is unparalleled in the known existence. This achievement is tied to the enormous communities of knowledge, made possible by language: leaving theological content aside, it is very much true that ``Εν αρχην ο λογος´´ (in the beginning was the word), and that in Western societies, this became particularly identified with the written word. There lies the challenge regarding modern age chatbots: they can ‘do’ language apparently as well as ourselves and there is a natural question of whether they can be considered intelligent, in the same way as we are or otherwise. Are humans uniquely intelligent? We consider this question in terms of the psychological literature on intelligence, evidence for intelligence in non-human animals, the role of written language in science and technology, progress with artificial intelligence, the history of intelligence testing (for both humans and machines), and the role of embodiment in intelligence. We think that it is increasingly difficult to consider humans uniquely intelligent. There are current limitations in chatbots, e.g., concerning perceptual and social awareness, but much attention is currently devoted to overcoming such limitations.
@@ 1 Statement of the Problem: are Chatbots Intelligent?
We are witnessing one of the most important and transformative technological revolutions in our history, the emergence of artificial intelligence (AI). AI may be compared with inventions of historic significance such as the printing press, steam engine, and the electronic computer, and even compared with these, the economic and scientific impact of AI is being felt during a much more concentrated period of time. A discussion with a modern chatbot, such as chat GPT (we refer to large language models as chatbots in a generic way), is undeniably human-like and several studies have been establishing that chatbots display human cognitive traits, including similar biases (binz2023using; suri2024decision), the recognition of emotional states (schaaff2023exploring) and empathy (roshanaei2024talk), and other mental processing similarities (mei2024turing; strachan2024testing). One effect of this revolution is existential: we, humans, often assume our place in the world is unique, by virtue of the uniqueness of our intelligence. Is human intelligence unique?
@@
A more practical question is whether chatbots will take our jobs. Together with being human-like, chatbots also encapsulate, in some sense, all human knowledge. Asking GPT-4 whether it will take my lecturer job, it is humble and reassuring: “As a university lecturer, you’re definitely in a role where expertise, creativity, and human connection are key. While AI, like ChatGPT, can help with things like answering questions, explaining concepts, or generating teaching materials, it’s not likely to replace the heart of your job.” However, increasingly in higher education, new policies are introduced to regulate the use of chatbots, given their dual ability to generate plausible coursework and to mark it with some competence (e.g., gultom2024use), questioning the role of humans in the process. Moreover, there are already indications that chatbots can carry out research roles (lehr2024chatgpt).
@@
The observation that technological advancements can lead to job losses is hardly new — there is a name for this, “technological unemployment” (peters2020beyond) — and there are some compelling examples, e.g., the reduction of the U.S. workforce employed in agriculture between 1900 (41%) and 2000 (2%, autor2014skills). Unsurprisingly, such concerns have been raised for chatbots too (Callum, 2016). The balancing perspective is that technological innovation also leads to new jobs (autor2015why; mincer2000technology) and over-productivity could be compensated in other ways (such as universal income). Additionally, some jobs require a human touch, do not depend on high levels of intelligence, or are even unnecessary (perhaps the result of vested interests). However, there is a unique role of chatbots in this debate in that they are not ‘just’ another technological innovation with potential to eliminate some jobs: in all cases of technological automation, the replacement concerned tasks that were either too laborious (e.g., agricultural tasks) or seen as beyond the competencies which we have considered central to our notion of human intelligence (e.g., multitudes of rote arithmetical operations in accounting). Instead, with chatbots, we are faced with an artificial agent which seems just like us, concerning some competencies which we have considered distinctly human.
@@
Overall, the consideration of whether chatbots can replace humans in various jobs is important and complements the existential question of whether chatbots are human-like. The present focus is intelligence. Can chatbots be considered intelligent? A difficulty in addressing this question is that our understanding of intelligence is itself incomplete. We consider some relevant ideas below. Briefly, on the one hand, it appears that traditional notions of intelligence can be covered by chatbots – and, moreover, a careful discussion reveals challenges to a uniqueness premise for human intelligence. On the other hand, the fundamental question is whether there is a special ingredient in human intelligence, which allows us to rise to apparent amazing intellectual accomplishments. What is it about our minds which, for example, has enabled us to invent quantum mechanics? Note, considering (something like) quantum mechanics an impressive intellectual accomplishment is not an anthropocentric bias: quantum mechanics has allowed surprising, transformative insights into the workings of nature, leading to unthinkable inventions such as semiconductors (and so modern computers). We think that the value of such accomplishments is manifest, independently of the fact that it was us who made them.
One possibility might be that there is no ‘special’ ingredient, rather human achievement is gradual and social, so we could possibly explain in detail why quantum mechanics and electronic computers would have been almost impossible to invent before industrial developments in the 1800s, without assuming that human intelligence is unique as a precondition.
@@ 2 What is Intelligence: Psychological Approaches
Attempts to define intelligence long predate modern psychology. In Western philosophy, intelligence has been traditionally connected with reason and thinking (gardner2011frames, Ch. 1). Aristotle associates reasoning particularly with the concept of λογος, a historic connection discussed in Section 4.
For Descartes, awareness of one’s own thoughts proved existence itself (cogito ergo sum). This tradition takes on a modern form in the work of Freud and Jung, who made introspection about the self and its symbolic life central to the emerging science of psychology.
There has also long been a desire for more empirical foundations — Hume, for example, was particularly sceptical of self-reported self-awareness as a scientific grounding. Mathematicians including Leibniz and Boole proposed ways to formalize and even automate reasoning, and Whitehead extended such symbolic approaches into broader theories of abstraction and meaning. These efforts eventually aligned intelligence with symbolic operations that could be both economically advantageous and systematically measured, which we’ll see became common practice in the later 1800s.
In psychology, by the late 1800s, figures such as Galton, Wundt, and William James had begun to turn these ambitions into methods, using reaction times, association tests, and other measures to make mental processes experimentally observable. By this point, Western psychology largely treated intelligence as a property of individual minds, though with different approaches to studying those minds.
@@
As certain kinds of intelligence could be tested, it became tempting to define intelligence in terms of what such tests could measure. Much of the psychological tradition regarding intelligence concerns psychometric theories, which aim to explain individual differences in test scores, on the basis of factors (e.g., see (sternberg2012intelligence)). The earliest proposal of this kind is due to spearman1904general, who argued for two factors, a general one (later associated with ‘mental energy’) and a specific one, comprised of competencies for particular tasks. thurstone1938primary, instead, identified seven “primary mental abilities” as the determinants of intelligence. These mental abilities include verbal fluency, spatial visualization, and perceptual speed. While originally spearman1904general and thurstone1938primary were interpreted as antithetical, this is not necessarily the case. One way to reconcile the two views is if the general vs. more specific intelligence factors are hierarchical, so that the latter informs specific aspects of the former. This hierarchical approach has been taken up by other intelligence psychologists, such as carroll1993human, who expanded both the specific and general abilities.
@@
This literature has recognised that some abilities reflect knowledge, while others processing power (in some general sense). cattell1943measurement introduced the distinction between fluid and crystallized abilities, with the former concerning performance that depends on prior knowledge minimally (e.g., problem-solving on a novel task) and conversely for the latter (e.g., vocabulary tasks). There is a converse relation between these two kinds of abilities and age, with fluid decreasing, while crystallized increasing, with age (e.g., ghisletta2012two).
@@
A more processing account of intelligence has been that it is characterized by the speed of information processing (deary2001reaction), sometimes in simple perceptual tasks, such as whether two lines are the same length (mackintosh1981measure) (but see (kiesling1988clocking)). There have been related proposals; for example, hunt1973individual considered the speed with which information is ’absorbed’ into memory and subsequently retrieved. Such approaches do not depend on extent of knowledge or other culture-specific skills, such as social skills.
@@
There have been other attempts to understand intelligence in more cognitive terms, e.g. in terms of working memory capacity (conway2013working), underpinned by high correlations between measures of working memory and reasoning ability (e.g., (kyllonen1990reasoning; ackerman2005working)). Interestingly, in cognitive science, working memory capacity limitations have been argued to enable learning in complex domains (newport1988constraints).
@@
Another processing idea is that intelligence relates to the ability to solve problems and achieve goals. sternberg1985beyond proposed that intelligence is characterized by three abilities: practical intelligence (environment interaction), creative intelligence (new problems and situations), and analytical intelligence (primarily problem-solving). Problem-solving ability has had a prominent role in other proposals (e.g., (danner2011measuring; hambrick2019problem)), as well as often being a component in most others (starting from (spearman1904general)). sternberg1985beyond’s theory is an example of approaches positing multiple types of intelligence. gardner2006multiple proposed a longer list, including musical, interpersonal, spatial, and linguistic intelligence, and salovey1990emotional introduced the idea of emotional intelligence as the ability to perceive and utilize emotions in cognition.
@@
Psychometric approaches to intelligence have demonstrated practical value. For example, they predict longevity (deary2001intelligence; deary2007intelligence), health outcomes (hart2003childhood), job performance (schmidt2004general; ericsson2016peak), and academic success (deary2007intelligence).
@@
Particular aspects of intelligence have been prioritised in intelligence testing: the approaches of (spearman1904general), binet1905methodes were translated into educational and cultural settings that emphasized the importance of written answers to prearranged test questions (rogele2020biographies). A preference for written demonstrations of intelligence, particularly in English-speaking educational systems, has profoundly affected the way artificial intelligence has been viewed and pursued, a relationship that we explore throughout the rest of this paper.
@@
Summing up, this is a complex and rich literature, and we only want to touch upon a handful of points, as relevant to this discussion. First, it is assumed that there are cognitive abilities that correlate with intelligence/the intelligences (corr2024personality). Second, some of these abilities concern processing characteristics, for example, processing speed or memory efficiency. Third, intelligence involves several aspects, e.g., problem-solving, linguistic/verbal, emotional, social, etc. Can these insights help us understand what, if anything, sets apart human intelligence compared to biological species and artificial agents? We can eliminate some possibilities. Artificial agents, including chatbots, exceed human capacity in essentially all processing characteristics (speed of processing, memory, information absorption and retrieval, problem solving ability, at least in specific categories), that is, it is unlikely that we can understand human vs. chatbot intelligence in terms of processing characteristics. Regarding types of intelligence, note first that non-human animals often demonstrate the competencies (at least to a degree) cited for humans too, such as spatial, emotional, musical abilities etc. (cf. gardner2006multiple).
Also, work such as Gardner’s argues for many types of intelligence, over and above linguistic and logical intelligence. There is certainly a question of whether there are types of human intelligence less possible for chatbots. As we shall see later, psychometric testing of human intelligence has focussed mostly on linguistic and logical intelligence, leading to challenges which are easily won by chatbots, both because of content and format (e.g., multiple choice tests, while cheap because of automated marking, are particularly easy for chatbots). Indeed, regarding psychometric testing, the linguistic competence of most chatbots is mostly equivalent to that of humans, even if the extent of understanding is unclear (mahowald2024dissociating). Regarding types of intelligence beyond linguistic and logical, we return to this issue when we discuss embodiment.
@@ 3 Animal Intelligence
Humans are not the fastest, toughest, most enduring etc. creatures, but we are distinguished for our cultural, technological, and scientific accomplishments – our achievement in these respects surpasses that of all other creatures. This has allowed us to consider ourselves the most intelligent species. Even if this is true in some sense, there is a question of whether other species are intelligent like us, but to a lesser extent, less intelligent in fundamental ways, or perhaps in some ways more intelligent than humans have realised.
@@
There are signs of intelligence throughout the animal kingdom (Bukart et al., 2017). For example, dogs have highly developed social awareness, including sensitivity to human body language, human gestures, and parts of language (hare2002domestication; kaminski2004word). Dogs are able to solve a variety of logical/ reasoning problems, such as a means end task (range2011dogs) or spatial puzzles (barrera2015effects). Theory of mind is the ability to attribute mental states (such as beliefs and desires) to others. It is typically considered an integral part of intelligence, especially social intelligence (navarro2022what). There have been arguments that dogs and other primates have a theory of mind (catala2017dogs; premack1978chimpanzee), though there is controversy (heyes1998theory). Chimpanzees have extensive tool-using abilities, including for food gathering/ hunting, grooming, and communication (koops2015chimpanzees). Chimpanzees can recognize abstract symbols (rumbaugh2003intelligence) and can innovate solutions to unfamiliar problems (bandini2020innovation). Similar evidence has accumulated for other species, e.g., corvids, who appear to have problem-solving abilities comparable to those of chimpanzees (gunturkün2023birds). Perhaps unsurprisingly, there is much evidence regarding musical intelligence in birds, e.g., relating to rhythm perception and processing (fitch2015four; patel2006musical) or the structural properties of birdsong (kawaji2024goal; zhang2023analogies). Further, there is some evidence of musical synchronisation amongst birds, though for duetting birds at most (hall2009review; mann2006antiphonal). Overall, bird musical ability does not appear to match the complexity of human music (lerdahl1983generative; levitin2006brain).
@@
What about, for example, the homing instinct of a pigeon or the way a leopard calculates the exact time and distance before it pounces onto prey? For example, pigeons have been known to successfully return home across distances exceeding 100km (schmidt1972homing). We would not call an individual pigeon intelligent for demonstrating its homing instinct, in the same way we would not call an individual human intelligent for learning to walk or ride a bicycle — such skills are considered default competencies that most members of a species are endowed with, as a result of evolutionary adaptation. There is intelligence at the species level, allowing for different competencies, developed through evolution (e.g., (jerison1988intelligence)). A similar case can be made regarding the remarkable ingenuity in, for example, the way ants or bees build their nests (winston1987biology). Note, linguistic competence could also be considered an instinct (e.g., pinker2015language). It is not the linguistic competence as such which is the crucial point, we think, but rather the fact that complex thought is typically (not always) communicated linguistically. Modern chatbots are clearly linguistically competent. Are they also capable of complex thought to the extent that we can consider them intelligent?
Self-interest is evident throughout animal behavior (for example, taking steps to avoid being eaten). Self-awareness has also been studied in animals, including the famous finding of gallup1970chimpanzees that chimpanzees, but not monkeys, learn to recognize a reflection in a mirror as their own. It is tempting to assume that humans are uniquely self-aware, hence more intelligent than other animals, but attempts to describe or demonstrate something distinctly self-aware about humans are empirically challenging (carruthers2009how). Whatever form mental states may take, humans can communicate limited parts of these experiences to one another, yet lack any comparable window on the mental life of animals; hence, we should not take absence of evidence as evidence of absence.
Symbol processing has long been emphasized as a traditional core of intelligence (e.g. whitehead1927symbolism). The recent ability of chatbots to manipulate linguistic symbols so fluently is disconcerting partly because, until now, this capacity seemed unique to humans. Yet, as torresmartinez2024embodied argues, non-human animals possess distinct semiotic systems that shape their own forms of cognition. He suggests that robot and AI design should attend to these alternative systems to avoid reproducing the specifically human constraints embedded in our symbolic models of thought.
@@
Beyond evolutionary intelligence, for some species, there is collective intelligence, the ability to achieve more than what an individual could accomplish, through cooperation. In the animal kingdom, for example, ants use pheromones to communicate about food sources or dangers and often work together in foraging tasks (holldobler1990ants); the way African wild dogs coordinate in hunting has sophisticated characteristics, including role differentiation and vocal communication (creel2002african; for an example regarding see rats schuster2002cooperative). However, no other species has been more successful in transcending what is possible for an individual than humans. fernbach2018knowledge coined the term ‘community of knowledge’, for the idea that even seemingly simple aspects of our lives are underwritten by massive knowledge contributions of which we are mostly ignorant. Our ability to form sophisticated communities of knowledge is fundamentally linked to the unique expressivity of human language (e.g., lakoff2003metaphors; pinker1995language) and the fact that our language has a written form (latour1992laboratory). Written language has enabled global communities of knowledge; arguably, it is the cornerstone of our civilization (goody1977domestication; concerning the interaction between climate, the durability of early writing mediums, such as papyrus, and the emergence of early civilizations see gaudet2014papyrus).
@@
Is there any evidence for art, science, technology amongst non-humans? Regarding art, some animals can learn to discriminate based on the aesthetics of pictures and show preference towards higher aesthetics (e.g., songbirds, for musical stimuli; watanabe2012animal). Some birds create and decorate elaborate structures to attract mates (borgia1985bower) and some animals can create pictorial patterns in the sand or using rudimentary tools, which resembles artistic expression (pele2021orangutan). However, such evidence is a far cry from demonstrations that non-human animals show human-like appreciation of these behaviours and their outputs, in the way we do for art (english2014painting). Regarding science, the scientific endeavour implies experimentation and exploration of testable hypotheses, to identify general principles and so generalize from known situations to novel ones. There is no evidence that animals can do science. However, there is some evidence that animals can experiment and adjust strategies, create artifacts, and develop rudimentary tools. For example, chimpanzees make tools to help them forage for ants or termites, through a process of successive improvement before the desired form is achieved, sometimes before actual tool use (indicating that they have an idea of the required form for a tool prior to use; boesch1990tool); and New Caledonian crows have been reported to construct hooked tools, to aid them in insect capture (hunt1973individual). Moreover, there is evidence of animals adapting behaviour from trial and error and changing circumstances (e.g., fiorito1992observational).
Such evidence indicates that animals are capable of technology and innovation, though it appears that their only path is through gradual trial and error, as opposed to the sudden realization which often accompanies human technological innovation (famous examples are penicillin and the discovery of the Benzene chemical structure, (fleming1929antibacterial; kekule1865constitution), as cited in rocke2015daydream).